### v1

# Achieving Complex Image Edits via Function Aggregation with Diffusion Models

Abstract
Diffusion models have demonstrated strong performance in generative tasks, making them ideal candidates for image editing. Recent studies highlight their ability to apply desired edits effectively by following textual instructions, yet two key challenges persist. First, these models struggle to apply multiple edits simultaneously, resulting in computational inefficiencies due to their reliance on sequential processing. Second, relying on textual prompts to determine the editing region can lead to unintended alterations in other parts of the image. In this work, we introduce FunEditor, an efficient diffusion model designed to learn atomic editing functions and perform complex edits by aggregating simpler functions. This approach enables complex editing tasks, such as object movement, by aggregating multiple functions and applying them simultaneously to specific areas. FunEditor is 5 to 24 times faster inference than existing methods on complex tasks like object movement. Our experiments demonstrate that FunEditor significantly outperforms recent baselines, including both inference-time optimization methods and fine-tuned models, across various metrics, such as image quality assessment (IQA) and object-background consistency.
Introduction
Diffusion models (DM) like DALLE-3 and Adobe Firefly have significantly advanced the field of image editing, marking a paradigm shift in how digital content is generated and refined. These models have revolutionized the creative process by allowing users to produce high-quality, photorealistic images from simple textual descriptions, a capability known as text-to-image (T2I) generation. Beyond generating images, diffusion models have also excelled in following text-based instructions to perform specific editing tasks, referred to as instruction-based editing. This approach makes image manipulation more intuitive and accessible by allowing users to guide the editing process with natural language commands (Brooks, Holynski, and Efros 2023; Sheynin et al. 2024; Geng et al. 2024). One prominent example of an instruction-based DM is InstructPix2Pix (IP2P) (Brooks, Holynski, and Efros 2023), which can edit images using textual instructions. EmuEdit (Sheynin et al. 2024) further improves the generalizability of IP2P by training on more diversified tasks. They found that training on multi-task datasets can enhance the models’ editing capabilities when handling challenging instructions.
Figure 1: (a) Results of applying basic functions, edge enhancement and object removal, on the corresponding masks. Best viewed when enlarged. (b) By applying each basic function simultaneously to different masks, we can perform complex edit functions. $\mathcal{A}$ is an aggregation function used for applying both functions in one shot.
Figure 2: Our approach is capable of composing multiple editing functions and applying them simultaneously. This enables it to perform complex edit functions such as object movement, object resizing, and object pasting in 4 steps. $f_{OR}$, $f_{EE}$, and $f_{HR}$ refer to object removal, edge enhancement, and harmonization functions, respectively. Each function is applied only to the specified mask region. To save space, source image I is omitted from the function arguments.
In recent years, the scope of instruction-based editing has expanded to include more complex tasks such as object movement (Wang et al. 2024; Mou et al. 2024a), subject composition (Yang et al. 2022), and virtual try-on (Chen et al. 2024b). In real-world scenarios, a single edit may involve multiple types of modifications to an image, resulting in a complex edit. For instance, the object movement task involves a combination of simpler editing functions: the model needs to fill in the missing region at the original object’s location with a natural background, while enhancement functions are often required for the moved object to ensure consistency with its new location. Recent studies (Guo and Lin 2024) reveal that instruction-based DMs, such as IP2P, struggle to faithfully follow editing instructions when multiple edits are required on an image, highlighting potential areas for improvement. Moreover, both IP2P and EmuEdit are limited to sequential editing, where each subsequent edit is performed on the results of the previous one. This approach increases the number of denoising steps during inference and is prone to error accumulation. To address this challenge, developing methods that can efficiently handle complex editing tasks is essential.
Previous works have attempted to tackle complex tasks either through inference-time optimizations (Mou et al. 2024b, a) or supervised fine-tuning (Wang et al. 2024; Winter et al. 2024). DiffEditor (Mou et al. 2024a) and DragonDiffusion (Mou et al. 2024b) rely on inference-time optimizations using energy-based loss functions, which involve an additional outer optimization loop and multiple calls to the UNet during inference, leading to high latency. Supervised fine-tuning approaches, such as AnyDoor (Chen et al. 2024b), address this issue by directly training a diffusion model capable of complex edits. However, collecting training data for all possible complex tasks is challenging, often requiring significant cost and human effort.
In this work, we propose FunEditor, an efficient diffusion model that performs complex edits through the aggregation of simpler functions. Under this formulation, FunEditor learns to effectively integrate multiple localized edit functions and apply them simultaneously. This is achieved by defining learnable task tokens during training and combining them during inference. For localization, each simple task requires a binary mask as input and performs cross-attention modification (Guo and Lin 2024; Simsar et al. 2023). Compared to prior editing models, FunEditor offers three unique advantages. First, during training, it only requires simple and atomic editing tasks such as object removal, harmonization, edge enhancement and etc. to learning new task tokens. During inference, by simply passing a sequence of these task tokens it applies multiple atomic functions simultaneously, requiring no energy functions. Figure 1 highlights the output of FunEditor on object movement and object resizing by aggregating object removal and edge enhancement. Second, FunEditor is data-efficient as we only require training on each atomic task individually, where training data is abundant or easy to collect. Making it the superior, low-cost option when compared to prior training-based approach like AnyDoor (Chen et al. 2024b). By training on atomic tasks it could generalize to combining them and achieve complex edits like object movement during inference. Third, FunEditor prioritizes localized editing where each simple edit function only updates a specified region in the input image. This leads to improved consistency. Furthermore, as FunEditor is compatible with pre-trained few-step diffusion models such as LCM (Luo et al. 2023), these atomic functions can be combined with pixel-space operation like copy and pasting to perform high-quality simultaneous edits in just 4 steps without the need of subject regeneration or extra optimization steps during inference as indicated in Figure 2.
To evaluate the effectiveness of FunEditor, we conducted extensive evaluations on object movement and object pasting as complex editing tasks using the COCOEE and ReS datasets. We compared FunEditor against a wide range of recently proposed editing baselines, including both inference-time optimization methods (Mou et al. 2024b) and fine-tuned models (Chen et al. 2024b). The results demonstrate that FunEditor significantly outperforms all baselines across various evaluation metrics, such as IQA and object-background consistency. Notably, FunEditor performs complex edits in just 4 steps, making it 5 to 24 times faster than the baselines in the object movement task. In addition to quantitative results, visual comparisons in Figures 5 and 6 further highlight FunEditor’s superiority in complex editing tasks.
Related Work
Image Editing with Diffusion Model.
The success of text-to-image (T2I) diffusion models in generating high-quality images has led to their widespread adoption in image editing (Epstein et al. 2023; Wang et al. 2024; Mou et al. 2024b, a; Alaluf et al. 2023; Hertz et al. 2023; Jeong et al. 2024; Patashnik et al. 2023). Training-free editing methods, such as Prompt-to-Prompt (Hertz et al. 2022), leverage these models by swapping cross-attention maps (CA-maps) of specific tokens during inference, enabling real-time editing without the need for fine-tuning. Plug-and-Play (Tumanyan et al. 2023) extends this approach by introducing spatial features and attention maps to guide the generation process and localize changes using masks. These methods exploit the generative capabilities of T2I diffusion models to apply edits and precisely target specific areas. Another line of work defines the editing function during inference using an energy function to update the noisy latent space (Mou et al. 2024a; Epstein et al. 2023; Shi et al. 2023; Mou et al. 2024b). This approach guides the editing direction and samples noise to align the latent space with the editing goal, as demonstrated in DragonDiffusion (Mou et al. 2024b) and DiffEditor (Shi et al. 2023). However, these techniques often require additional steps, such as optimization and image inversion using methods like DDIM (Dhariwal and Nichol 2021), which can reduce efficiency.
Training-based Editing Model.
As the editing functions become more complex, the T2I diffusion model may show inconsistencies and poor performance. This is primarily because the model has not seen similar examples during training. Training-based approaches address this by using large editing datasets (Zhang et al. 2024a, b; Hui et al. 2024), which include pairs of source and edited images along with the corresponding edit instructions, to fine-tune a T2I diffusion model specifically for editing tasks. InstructPix2Pix (IP2P) (Brooks, Holynski, and Efros 2023) leverages GPT-3 (Brown et al. 2020) and Prompt-to-Prompt (Hertz et al. 2022) to generate synthetic editing datasets. To enable the source image to be passed to the UNet as an additional input, they added a new channel to the input convolutional layer, allowing the model to preserve details in unintended areas. Building on the success of InstructPix2Pix (IP2P), subsequent works, such as MagicBrush (Zhang et al. 2024a), have manually annotated images to create higher-quality datasets. By utilizing the same training procedure to fine-tune their editing models, these approaches have demonstrated improved performance (Zhang et al. 2024b; Hui et al. 2024). However, the main shortcoming of IP2P-based editing models is the occurrence of unintended changes to other parts of the source image (Simsar et al. 2023; Guo and Lin 2024). To address this issue, FoI (Guo and Lin 2024) introduced an inference technique that modulates attention and masks the sampled noise for the text-conditioned output, thereby limiting changes only to the specific area of the edit.
Multi-task Diffusion Model.
Unifying multiple tasks within a single model has been an intriguing research direction in the literature. InstructDiffusion (Geng et al. 2024) investigates the potential benefits of joint training across various tasks, including keypoint detection, segmentation, image enhancement, and editing. Their results suggest that multi-task training enhances the model’s performance across these diverse tasks. Similarly, EmuEdit (Sheynin et al. 2024) introduces a unified diffusion model capable of performing multiple editing and computer vision tasks, such as object removal, global editing, and object detection. EmuEdit uses task embeddings as conditions to switch between tasks, effectively allowing the model to adapt its behavior to the specific task at hand. While these approaches show impressive results on individual tasks, they are limited by their inability to apply two tasks simultaneously, requiring multiple user interactions to achieve complex edits. In this work, we address this limitation by bridging the gap between multi-task training and the ability to perform multiple editing tasks in a single operation, enabling efficient complex editing.
Method
In this section, we explain how complex localized editing tasks could be formulated into a problem of function aggregation. This formulation allow us to achieve simultaneous edits during inference in few-steps. Then, we showcase how FunEditor solve this problem by a carefully designed training framework. Figure 3 provides an overview of our proposed pipeline for both training and inference.
Figure 3: Overview of our proposed training and inference pipeline. During the basic task training phase (a) the diffusion model learns to perform various simple tasks based on the provided task tokens and masks. During inference (b), we could implement complex edit functions by combining multiple task masks and tokens.
Editing as Function Aggregation
Let $f$ be a localized editing function to a certain region of the source image $I$, indicated by a binary mask $m$. $f(I,m)$ denotes the edited image. We then define an aggregation function $F$ as:
$F=\mathcal{A}(f_{1}(I,m_{1}),f_{2}(I,m_{2}),\dots,f_{n}(I,m_{n})).$ (1)
In essence, $F$ represents aggregating different $f$ and applying them simultaneously to multiple regions of the source image in one shot. The main advantage of such formulation is that it enables the divide-and-conquer of complex tasks (see Figure 1). Given that it is challenging to collect sufficient high-quality training data to directly learn $F$, our goal is to learn each $f$ individually during training and aggregate their effects during inference. In FunEditor, each basic function $f_{i}$ is represented by a trainable task token $\mathbf{T}_{i}$. For example, providing the token <Task-OR> to the UNet as input would enable the task of object removal, as indicated by Figure 3(a). Although EmuEdit also define task embedding vectors that are of similar purposes, the main difference is that we directly add the task tokens to the vocabulary of the prompt tokenizer. This design offers two advantages compared to EmuEdit. First, it allows FunEditor to freely combining tasks by passing the corresponding task tokens as prompts to the UNet. To disable tasks, we introduce another special token named <SKIP-TASK>, which ensures a consistent prompt length during training and inference, improving the overall generalizability. Second, task tokens pave the way towards localized, simultaneous editing during inference. To achieve this, we simply set the desired task tokens in the input prompt and keep the remaining ones as <SKIP-TASK>. Similar to instruction-based editing (Simsar et al. 2023; Guo and Lin 2024), our goal is to apply each editing function $f_{i}$ to its corresponding mask $m_{i}$ to maintain control over the editing region and prevent leakage into other areas.
Since the UNet processes the task tokens via cross-attentions (CA), it is natural to inject localization information via the same medium. Therefore, we employ a cross-attention (CA) masking technique to ensure each task $f_{i}$ does not affect the unmasked area $1-m_{i}$. During both training and inference, we apply this CA masking technique to all CA layers. For each CA layer, we interpolate the mask into its resolution and set the values of $QK^{T}$ of $\mathbf{T}_{i}$ token corresponding to the unmasked area $1-m_{i}$ to its minimum value. Mathematically represented as:
$QK^{T}[(1-m_{i}),\mathbf{T}_{i}]=\min(QK^{T}).$ (2)
Figure 4 highlights the importance of CA masking for localized editing. Notably without CA masking the model could attend to other regions of the source image and generates undesired changes outside of the masked region.
Figure 4: Harmonization without cross-attention masking affects the entire image (a). While with masking, edits are confined to the masked region, (b), preventing changes to unmasked areas. Mask is indicated by the top right mini-figure.
Figure 5: Qualitative comparison between our approach and baseline methods for object repositioning within an image, demonstrating the superior performance of our method. To move an object, FunEditor composes object removal and edge enhancement functions.
Figure 6: Visual comparison between our method and baseline methods for object pasting from a reference image into a target image. FunEditor applies harmonization and edge enhancement functions to seamlessly paste an object into another image.
FunEditor Training Strategy
The training procedure for FunEditor builds upon the IP2P architecture. In IP2P, a source image $I_{src}$, a target image $I_{trg}$, and a text instruction $T$ are provided, with noise $\epsilon$ added to $I_{trg}$ to create a noisy input $z_{t}$. The UNet $\epsilon_{\theta}$ is trained to predict $\epsilon$ using the following objective:
$\min_{\theta}\mathbb{E}_{y,\epsilon,t}\left[||\epsilon-\epsilon_{\theta}(z_{t},t,I_{src},T)||^{2}_{2}\right],$ (3)
where $y=(I_{src},I_{trg},T)$ represents a sample from the IP2P editing dataset. In FunEditor, instead of using a text instruction $T$, we incorporate new task tokens into the CLIP text tokenizer (Radford et al. 2021) to represent the desired editing function. Additionally, we use masks corresponding to each editing function within the cross-attention. Thus, our training sample takes the form $(I_{src},\mathbf{T},m,I_{trg})$, where $I_{src}$ is the source image, $\mathbf{T}$ is the task token, $m$ is the mask, and $I_{trg}$ is the target image.
The idea of task embeddings provide the foundation for enabling simultaneous edits in FunEditor. EmuEdit defines learnable task embedding vectors, which are provided to the UNet as task-selectors via cross-attention interaction and the timestep embedding layers. In contrast, we incorporate the task embeddings via cross-attention to have more control on their spatial effects on the editing regions. The following equation defines our objective for training FunEditor, where we learn an embedding vector for each task token that represents a basic edit:
$\min_{\theta,v_{1},\ldots,v_{k}}\mathbb{E}_{\hat{y},\epsilon,t}[||\epsilon-\epsilon_{\theta}(z_{t},t,I_{src},m,\mathbf{T})||^{2}_{2}].$ (4)
Experimentally we find that setting the number of task tokens in the prompt equal to the number of supported simple tasks generates the best results. During training we randomly activate one task token and keep the remaining tokens as <SKIP-TASK>. Since the task token ordering is not important, we also randomly shuffle the task tokens in every training step to prevent the model from memorizing their positions.
Experiments
In this section, we evaluate the effectiveness of our proposed method by evaluating it on to two complex image editing tasks: Object Movement ($F_{OM}$) and Object Pasting ($F_{OP}$). Building on the formulation introduced in the previous section, we define each complex function as an aggregation of the following basic functions:
$F_{OM}(I,M_{src},M_{trg})=\mathcal{A}\big{(}f_{OR}(I,M_{src}),f_{EE}(I,M_{trg})\big{)},$
$F_{OP}(I,M_{trg})=\mathcal{A}\big{(}f_{HR}(I,M_{trg}),f_{EE}(I,M_{trg})\big{)}.$
In these formulations, $f_{OR}$, $f_{EE}$, and $f_{HR}$ represent the basic functions for object removal, edge enhancement, and harmonization, respectively. The masks $M_{src}$ and $M_{trg}$ correspond to the source object and the target location within the image.
Our method is compared to baseline approaches with respect to both quality and efficiency. The subsequent subsections will detail the experimental setup, including the datasets used and evaluation metrics. We will then present the results of our experiments, comparing the performance of our method with that of baseline techniques across the defined tasks.
Method #Steps NFEs Avg. Latency std
SelfGuidance 50 100 11 0.6
DragonDiffusion 50 160 23 0.7
DiffEditor 50 176 24 0.4
Paint-by-Example 50 50 5 0.6
AnyDoor 50 50 12 0.4
FunEditor (ours) 4 4 1 0.4
Table 1: Efficiency is compared in terms of the number of function evaluations (NFEs) and latency (seconds). Latency is measured as the average wall-clock time for editing one image over 10 runs on a single Nvidia V100 (32GB) GPU.
Experimental Settings
Implementation Details
We fine-tuned our model using 4 Nvidia V100 (32GB) GPUs, with the IP2P UNet (Brooks, Holynski, and Efros 2023) serving as our diffusion backbone. The AdamW optimizer (Loshchilov and Hutter 2017) was employed with a learning rate of $5\times 10^{-5}$ and a batch size of 4. During inference, we distilled the trained model into 4 steps using SD1.5 LCM-LoRA from HuggingFace [^1] https://huggingface.co/latent-consistency/lcm-lora-sdv1-5. For the baselines, we used SDv1.5 (Rombach et al. 2022) from HuggingFace [^2] https://huggingface.co/runwayml/stable-diffusion-v1-5, adhering to all default inference hyperparameters. As a preprocessing step for object removal, we dilated the mask with a kernel size of 20 to prevent any pixel leakage into the inpainted region.
Datasets
The COCOEE dataset is an editing benchmark compiled by Yang et al. (2022), featuring 3,500 images manually selected from the MSCOCO (Microsoft Common Objects in Context) (Lin et al. 2014) validation set. A human operator used the Segment Anything model (Kirillov et al. 2023) to extract segments from a random subset of COCOEE, identifying suitable objects and assigning diff vectors to specify their new locations, resulting in a benchmark of 100 images with corresponding masks and diff vectors for object movement tasks. In addition to our annotated COCOEE dataset, Wang et al. (2024) released the ReS dataset, comprising 100 pairs of real-world images that present challenging cases of object movement. The details of the training datasets for atomic edits can be found in the appendix.
Method IQA Object Consistency Background Consistency Semantic Consistency
TOPIQ $\uparrow$ MUSIQ $\uparrow$ LIQE $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ CLIP-I2I $\uparrow$
SelfGuidance 0.554 65.91 3.90 0.083 22.77 0.259 17.86 0.897
DragonDiffusion 0.571 68.87 4.27 0.034 28.59 0.098 23.99 0.965
DiffEditor 0.579 69.09 4.27 0.036 28.49 0.094 24.23 0.967
Paint-by-Example 0.595 69.57 4.31 0.066 23.12 0.089 25.23 0.937
AnyDoor 0.558 68.37 4.12 0.053 25.43 0.122 23.55 0.954
FunEditor (ours) 0.611 70.39 4.36 0.017 34.27 0.066 26.54 0.969
Table 2: Quantitative evaluation of our approach compared to the baselines on the object movement task using the COCOEE dataset. To move an object, FunEditor encompasses object removal and edge enhancement functions.
Metrics
In our evaluations, we utilize four groups of metrics: overall image quality, object consistency, background consistency, and semantic consistency. For Image Quality Assessment (IQA), we assess the perceptual quality of the images using TOPIQ (Chen et al. 2024a), MUSIQ (Ke et al. 2021), and LIQE (Zhang et al. 2023). Object Consistency is measured by comparing the similarity of the moved object to its original counterpart in the source image, using the LPIPS (Zhang et al. 2018) and PSNR metrics. Background Consistency is evaluated by assessing the similarity between the background in the edited image and the original image, using the same metrics as for Object Consistency. Finally, Semantic Consistency is determined by evaluating the similarity between the semantics of the source and edited images through CLIP image encoder, which measures image-to-image similarity based on CLIP embeddings (Radford et al. 2021).
Baselines
We evaluate our approach against both training-based and training-free methods on complex editing tasks. Given a limited number of open-source training-based model specifically for complex tasks (Winter et al. 2024; Wang et al. 2024), we include training-free approaches to comprehensively assess the quantitative performance of our method. Our baselines for the complex editing tasks are as follow:
  • Paint-by-Example (Yang et al. 2022): This study explores a training-based approach for reference-guided image editing, training a model to fill a designated mask area with a reference object.
  • AnyDoor (Chen et al. 2024b): A training-based method that enables teleportation of an object into a user-specified location and shape, leveraging object identity information to accurately reconstruct the object in the intended form.
  • SelfGuidance (Epstein et al. 2023): A training-free method that utilizes energy functions on attention maps to optimize the noisy latent representation during inference, allowing for edits such as object movement.
  • DragonDiffusion (Mou et al. 2024b): A training-free approach that employs an energy function with two components: one for applying changes within the target mask and another for preserving the unmasked area.
  • DiffEditor (Mou et al. 2024a): Similar to DragonDiffusion, this method enhances the quality of edits by optimizing an energy function, incorporating score-based gradient guidance and time interval strategies into the diffusion noise sampling process.
Method IQA Object Consistency Background Consistency Semantic Consistency
TOPIQ $\uparrow$ MUSIQ $\uparrow$ LIQE $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ CLIP-I2I $\uparrow$
SelfGuidance 0.586 69.41 3.61 0.064 24.21 0.273 17.92 0.869
DragonDiffusion 0.690 74.95 4.72 0.030 29.68 0.083 25.38 0.934
DiffEditor 0.691 74.94 4.73 0.032 29.59 0.083 25.44 0.933
Paint-by-Example 0.717 74.94 4.75 0.052 24.20 0.076 26.27 0.907
AnyDoor 0.679 73.68 4.51 0.045 25.89 0.106 23.48 0.923
FunEditor (ours) 0.702 75.00 4.75 0.015 35.61 0.063 26.48 0.942
Table 3: Quantitative evaluation of our approach compared to the baselines on the object movement task was conducted using the ReS dataset. To move an object, FunEditor encompasses object removal and edge enhancement functions.
Experimental Results
Efficiency
Table 1 demonstrates that FunEditor, using only $4$ steps, requires $172$ fewer NFEs and is $23$ seconds faster than the state-of-the-art method for object movement, DiffEditor (Mou et al. 2024a), on the COCOEE dataset. SelfGuidance, DragonDiffusion, and DiffEditor, rely on optimizing energy functions to guide complex edits during inference. These method requires multiple calls to the UNet, as it must perform both inpainting and blending of the object with its surrounding environment, leading to a more time-consuming process. Comparing to training-based approaches, our model is $12$ times faster than the training-based baseline AnyDoor (Chen et al. 2024b) and $5$ times faster than Paint-by-Example (Yang et al. 2022). These methods regenerate the object to fit the target position, which also demands a higher number of steps to achieve the desired result. In contrast, FunEditor achieves superior results in fewer steps by employing simpler enhancement functions, such as edge enhancement, and pixel operations, without the need for fully regenerating the object. Instead, these functions refine the composition and harmonization of the pasted object in pixel space, enabling us to achieve complex edits more efficiently. It should be noted that the studies that used synthetic data generation for object movement (Winter et al. 2024; Wang et al. 2024) did not open-source their models, therefore we could not use them as baselines to assess their latency in a multi-turn fashion.
Image Quality
We evaluate the performance of FunEditor through quantitative and qualitative experiments on object movement and object pasting. Tables 2 and 3 presents our quantitative comparison with the baselines on the object movement task on COCOEE and ReS datasets. Our model achieves higher quality across almost all IQA metrics compared to the baselines, improving the accuracy of object movement while preserving the details of both the object and the background. By aggregating atomic functions, FunEditor successfully removes objects, enhancing semantic consistency and overall image quality compared to optimization-based approaches that rely on energy functions. Additionally, since we use the original object in the pixel space and apply enhancement functions such as edge enhancement and harmonization, we can faithfully move or paste the object without altering its appearance, unlike the baselines that regenerate the object and modify its appearance.
Figure 5 and 6 provide qualitative comparisons between our method and the baselines for object movement and object pasting. As shown in Figure 5, the baselines face two primary issues: unsuccessful object removal and unsuccessful object pasting. In contrast, our method seamlessly repositions and blends the object into the destination while effectively removing the original object. Supporting the quantitative results in Table 2, FunEditor faithfully preserves the background, applying changes only to the intended editing area. For example, SelfGuidance and DiffEditor alter the shape of the floor lamp in the first image. In the case of object pasting, as shown in Figure 6, regenerating the object can lead to issues such as unfaithful or incomplete regeneration like the first example.
Conclusion
In this paper, we introduced FunEditor, an efficient editing approach that utilizes function aggregation to address complex editing tasks such as object movement. FunEditor offers several key advantages: it requires only publicly available datasets for atomic tasks, enables the simultaneous execution of multiple functions to achieve complex edits, and is compatible with few-step diffusion models. Our approach not only outperforms baselines specifically designed for complex tasks like object movement and object pasting in terms of quality but also offers superior efficiency. Additionally, we evaluated FunEditor across a wide range of numerical metrics and visual quality, comparing it against both training-based and training-free baselines.

### v2

@@
-Achieving Complex Image Edits via Function Aggregation with Diffusion Models
+FunEditor: Achieving Complex Image Edits via Function Aggregation with Diffusion Models
 Abstract
-Diffusion models have demonstrated strong performance in generative tasks, making them ideal candidates for image editing.
-Recent studies highlight their ability to apply desired edits effectively by following textual instructions, yet two key challenges persist.
+Diffusion models have demonstrated outstanding performance in generative tasks, making them ideal candidates for image editing.
+Recent studies highlight their ability to apply desired edits effectively by following textual instructions, yet with two key challenges remaining.
 First, these models struggle to apply multiple edits simultaneously, resulting in computational inefficiencies due to their reliance on sequential processing.
-Second, relying on textual prompts to determine the editing region can lead to unintended alterations in other parts of the image.
-In this work, we introduce FunEditor, an efficient diffusion model designed to learn atomic editing functions and perform complex edits by aggregating simpler functions.
+Second, relying on textual prompts to determine the editing region can lead to unintended alterations to the image.
+We introduce FunEditor, an efficient diffusion model designed to learn atomic editing functions and perform complex edits by aggregating simpler functions.
 This approach enables complex editing tasks, such as object movement, by aggregating multiple functions and applying them simultaneously to specific areas.
-FunEditor is 5 to 24 times faster inference than existing methods on complex tasks like object movement.
-Our experiments demonstrate that FunEditor significantly outperforms recent baselines, including both inference-time optimization methods and fine-tuned models, across various metrics, such as image quality assessment (IQA) and object-background consistency.
+Our experiments demonstrate that FunEditor significantly outperforms recent inference-time optimization methods and fine-tuned models, either quantitatively across various metrics or through visual comparisons or both, on complex tasks like object movement and object pasting.
+In the meantime, with only 4 steps of inference, FunEditor achieves 5–24$\times$ inference speedups over existing popular methods.
+The code is available at: mhmdsmdi.github.io/funeditor/.
 Introduction
@@
 They found that training on multi-task datasets can enhance the models’ editing capabilities when handling challenging instructions.
-Figure 1: (a) Results of applying basic functions, edge enhancement and object removal, on the corresponding masks.
-Best viewed when enlarged.
-(b) By applying each basic function simultaneously to different masks, we can perform complex edit functions.
-$\mathcal{A}$ is an aggregation function used for applying both functions in one shot.
+Figure 1: (a) Results of applying the two basic functions—Edge Enhancement (center) and Object Removal (right)—using their respective masks on the input image Best viewed when enlarged.
+(b) Demonstration of function aggregation using the proposed method.
+By simultaneously applying Object Removal and Edge Enhancement on different masks, complex edits such as object shrinking (middle) and object movement (right) can be achieved.
+$\mathcal{A}$ represents the operation of function aggregation.
 Figure 2: Our approach is capable of composing multiple editing functions and applying them simultaneously.
@@
 Compared to prior editing models, FunEditor offers three unique advantages.
-First, during training, it only requires simple and atomic editing tasks such as object removal, harmonization, edge enhancement and etc.
-to learning new task tokens.
-During inference, by simply passing a sequence of these task tokens it applies multiple atomic functions simultaneously, requiring no energy functions.
-Figure 1 highlights the output of FunEditor on object movement and object resizing by aggregating object removal and edge enhancement.
-Second, FunEditor is data-efficient as we only require training on each atomic task individually, where training data is abundant or easy to collect.
-Making it the superior, low-cost option when compared to prior training-based approach like AnyDoor (Chen et al.
+First, during training, it only requires simple and atomic editing tasks such as object removal, harmonization, edge enhancement, etc., to learn new task tokens.
+During inference, by passing a sequence of learned task tokens into diffusion model, FunEditor can apply multiple functions simultaneously to accomplish a complex edit, requiring no energy guidance during inference.
+Figure 1 highlights the output of FunEditor for object movement and object resizing by aggregating object removal and edge enhancement.
+Second, FunEditor is data-efficient as we only require training on each atomic task individually, for which training data is either abundant or easy to collect, making it a superior, low-cost option when compared to prior training-based approach like AnyDoor (Chen et al.
 2024b).
-By training on atomic tasks it could generalize to combining them and achieve complex edits like object movement during inference.
-Third, FunEditor prioritizes localized editing where each simple edit function only updates a specified region in the input image.
-This leads to improved consistency.
-Furthermore, as FunEditor is compatible with pre-trained few-step diffusion models such as LCM (Luo et al.
-2023), these atomic functions can be combined with pixel-space operation like copy and pasting to perform high-quality simultaneous edits in just 4 steps without the need of subject regeneration or extra optimization steps during inference as indicated in Figure 2.
+Third, FunEditor prioritizes localized editing where each simple edit function only applies to a specified region in the input image, leading to improved consistency.
+Lastly, as FunEditor is compatible with pre-trained few-step diffusion models such as LCM (Luo et al.
+2023), these atomic functions can be combined with pixel-space operation like copy and pasting to perform high-quality simultaneous edits in only 4 steps without the need of subject regeneration or extra optimization steps during inference as indicated in Figure 2.
 To evaluate the effectiveness of FunEditor, we conducted extensive evaluations on object movement and object pasting as complex editing tasks using the COCOEE and ReS datasets.
@@
 The results demonstrate that FunEditor significantly outperforms all baselines across various evaluation metrics, such as IQA and object-background consistency.
-Notably, FunEditor performs complex edits in just 4 steps, making it 5 to 24 times faster than the baselines in the object movement task.
+Notably, FunEditor performs complex edits in just 4 steps, making it 5–24$\times$ faster than the baselines in the object movement task.
 In addition to quantitative results, visual comparisons in Figures 5 and 6 further highlight FunEditor’s superiority in complex editing tasks.
@@
 The UNet $\epsilon_{\theta}$ is trained to predict $\epsilon$ using the following objective:
-$\min_{\theta}\mathbb{E}_{y,\epsilon,t}\left[||\epsilon-\epsilon_{\theta}(z_{t},t,I_{src},T)||^{2}_{2}\right],$ (3)
+$\min_{\theta}\mathbb{E}_{y,\epsilon,t}\left[||\epsilon-\epsilon_{\theta}(z_{t},t,I_{src},I_{trg},T)||^{2}_{2}\right],$ (3)
 where $y=(I_{src},I_{trg},T)$ represents a sample from the IP2P editing dataset.
@@
 Datasets
-The COCOEE dataset is an editing benchmark compiled by Yang et al.
-(2022), featuring 3,500 images manually selected from the MSCOCO (Microsoft Common Objects in Context) (Lin et al.
+The COCOEE dataset, compiled by Yang et al.
+(2022), features 3,500 images manually selected from the MSCOCO (Microsoft Common Objects in Context) (Lin et al.
 2014) validation set.
 A human operator used the Segment Anything model (Kirillov et al.
-2023) to extract segments from a random subset of COCOEE, identifying suitable objects and assigning diff vectors to specify their new locations, resulting in a benchmark of 100 images with corresponding masks and diff vectors for object movement tasks.
-In addition to our annotated COCOEE dataset, Wang et al.
-(2024) released the ReS dataset, comprising 100 pairs of real-world images that present challenging cases of object movement.
+2023) to extract segments and assign diff vectors, resulting in a benchmark of 100 images with corresponding masks and diff vectors for object movement tasks.
+Wang et al.
+(2024) also released the ReS dataset, comprising 100 pairs of real-world images that present challenging cases of object movement.
 The details of the training datasets for atomic edits can be found in the appendix.
@@
 2022).
-These methods regenerate the object to fit the target position, which also demands a higher number of steps to achieve the desired result.
+These methods regenerate the object to fit the target position demanding a higher number of steps to achieve the desired result.
 In contrast, FunEditor achieves superior results in fewer steps by employing simpler enhancement functions, such as edge enhancement, and pixel operations, without the need for fully regenerating the object.